# Learning policies for resource allocation in business processes

### v2

@@ 6 Conclusion
This paper presented two learning methods for resource allocation in business processes. The first approach is based on deep reinforcement learning (DRL), where an agent learns what assignments lead to a low cycle time by interacting with the environment. The second method is a score-based value function approximation (SVFA) approach, where a score function is learned that ranks assignments based on a set of curated features derived from the current process circumstances. We evaluate our proposed methods by benchmarking them against three traditional heuristics and the two most relevant existing BPO methods. We evaluate the methods on six business processes, also known as scenarios, representing archetypal process patterns and three realistically sized business processes that are a combination of these six scenarios.
The evaluation results show that our learning methods outperform or are competitive with the benchmarks in five out of six scenarios. Only in the scenario with parallelism did the First In, First Out heuristic outperform our methods because it prioritizes the activity that leads to case completion. Our methods and other benchmark methods do not explicitly consider information related to the execution sequence of cases, resulting in lower performance. One of the benchmarks from the literature was competitive with our methods in the scenario with high utilization, but neither of these existing methods could find a competitive policy for the other scenarios. These results showed that our learning-based methods are suitable for learning policies for process patterns commonly seen in business processes.
@@
The second part of the evaluation was performed on business processes created by combining the six scenarios in three different configurations: in sequence, in sequence with reversed order, and in parallel. While some heuristics were competitive or outperformed our methods in the scenarios, the heuristics exhibited limited performance when all process patterns from the scenarios were combined. Our proposed DRL method outperformed all methods in all three composite business processes and outperformed the best-performing benchmark by, on average, 13.1%. DRL automatically learns important features from the current state of the process, independent of the process characteristics, making it adaptive to different business process patterns and flows. Our DRL-based method is scalable and suitable for learning resource allocation policies in business processes. While SVFA was, on average, the best-performing method in the scenarios, it could not find a competitive policy in the composite business processes compared with DRL.
In future work, the current methods can be improved in several ways. One research direction is to consider other or multiple optimization objectives. While this work only focused on minimizing the cycle time, other objectives can be a more fitting choice depending on the setting. Furthermore, we trained and tested each model in the same environment. However, in a real-world process, resources may, for example, enter and leave the process, which current methods do not account for. Therefore, handling changes in the business process model is a relevant extension of our work. The process mining field has shown that the execution sequence of a running case, or prefix, holds valuable information but has not been used in any BPO approach so far. The scenario with the parallelism structure showed that the prefix can significantly improve performance, and we will investigate how we can use the prefix in our resource allocation algorithms. Furthermore, one limitation of DRL is that it is difficult for the agent to learn to what extent what action contributes to a specific reward realized many actions later. This behavior is caused by the design of the reward function. While it worked well in learning policies for the business processes considered in this paper, it may also lead to undesired behavior, as seen in the N-network. In future work, we aim to address the bias induced by the reward function.
@@
Reproducibility. The source code to reproduce the experiments can be found in our Github repository.

### v3

@@ 6 Conclusion
This paper introduced two learning methods for resource allocation in business processes: deep reinforcement learning (DRL) and score-based value function approximation (SVFA). We evaluated these methods on six archetypal business processes, or scenarios, and three composite processes, which are a combination of these scenarios in sequence, sequence reversed, and parallel.
Under a constant arrival rate, our methods outperformed or matched the benchmarks in five out of six scenarios. In the scenario with parallelism, our methods could not compete with the FIFO heuristic. The FIFO policy is a suitable and computationally efficient policy for resource allocation and its main strengths lie in prioritizing the quick completion of individual cases. This policy is especially effective in processes where choosing which activity to complete is more important than which resource to assign to it (i.e., slow downstream and parallelism scenario). Furthermore, our learning-based methods do not consider the prefix of a case, which could significantly improve the policy for the parallelism scenario. However, the disadvantage of FIFO, and heuristics in general, is that it can force sub-optimal assignments, while learning-based methods can adapt to different process states. The scenario results showed that our learning-based methods are suitable for learning policies for process patterns commonly seen in business processes.
@@
In the composite processes, our DRL method achieved a, on average, 12.7% lower cycle time compared to the best benchmark. Our SVFA method could not find a competitive policy and performed slightly worse than the best-performing benchmark. While some heuristics exhibited good performance in some scenarios, they could not handle the complexity of all scenarios combined. On the other hand, DRL automatically learns important features from the current state of the process, independent of the process characteristics, making it adaptive to different business process patterns and flows.
To investigate our methods further, we analyzed their performance when using an arrival pattern that adapts the arrival rate over time. The results showed that DRL and SVFA performed best in three and two scenarios, respectively. However, in the composite processes, our methods could not find a competitive policy. Adding a feature to DRL to represent the arrival pattern significantly improved the performance in the parallel scenario and the composite processes. After introducing this feature, DRL outperformed or matched the performance of the best-performing method in two out of three composite processes.
In future work, the current methods can be improved and extended in several ways. First, in this paper, we only considered a single objective, but other or multiple objectives, such as case outcome, can be more fitting depending on the process. Second, as shown in the results, our proposed methods cannot sufficiently handle an arrival pattern. Our methods were designed for constant arrival rates, and while adding a simple feature representing the arrival pattern significantly reduced the cycle time, more research is needed to handle more complex dynamics, such as seasonality in the process. Furthermore, additional scenarios should be researched to further investigate the adaptability of our methods. While we have already applied our DRL method to real-world business processes [33], we have not yet studied changes in the process over time, such as adding new employees. Third, process mining research has shown that the execution history of a case, or prefix, contains valuable information about future states. However, in current BPO methods, this information is not used and can be exploited to improve the optimization process. Lastly, in the N-network, the reward function encouraged the agent to focus on completing one of the two activities, resulting in high rewards, but also high cycle times. In future work, we aim to address the bias induced by the reward function and investigate a reward function that is more representative of the objective and does not require reward engineering.
@@
Reproducibility. The source code to reproduce the experiments can be found in our Github repository: http://www.github.com/jeroenmiddelhuis/
LearningResourceAllocation